Six measurements were made of the time in seconds for 20 oscillations of a mass hung from a spiral spring as follows: 20.1, 20.2, 20.3, 20.1, 20.0, 20.2. The mean value of the period of oscillation is
1
Expert's answer
2020-05-18T10:03:02-0400

Find the average value


"\frac {20.1+20.2+20.3+20.1+20.0+20.2}{6}=20.15"

Since the result corresponds to 20 oscillations, the period of one oscillation is:


"\frac {20.15}{20}=1.0075"